OpenStudy (anonymous):

Calculate the limit : lim (x->0-) ln(e^x)/(e^1/x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{-}} x^2/e = 0\]

OpenStudy (anonymous):

since\[ \ln (e^x) = x\]

OpenStudy (anonymous):

No, it's e^(1/x) not (e^1)/x ... my mistake...

OpenStudy (anonymous):

\[-\infty\]

myininaya (myininaya):

\[\frac{lne^\frac{1}{x}}{e^\frac{1}{x}}?\]

OpenStudy (anonymous):

Imagine it as so : if f(x) = e^x , then calculate : lim (x->0) lnf(x)/f(1/x)

OpenStudy (anonymous):

The solution is -inf but I need a more detailed answer please

myininaya (myininaya):

\[\frac{\ln(e^x)}{e^\frac{1}{x}}\]?

OpenStudy (anonymous):

Yes myininaya

OpenStudy (anonymous):

you need to use l'hospital's rule since you get a 0/0 indeterminate

OpenStudy (anonymous):

first you can simplify first to lim x-> 0- x/e^(1/x)

OpenStudy (anonymous):

then the derivative of the numerator is 1

OpenStudy (anonymous):

and the derivative of the denominator is e^(1/x) lnx

OpenStudy (anonymous):

Believe it or not this was in my high school exams and it gave you 5/100 points... Greece rocks...

OpenStudy (anonymous):

so you should get as a result lim x->0- 1/[e^(1/x) lnx]

OpenStudy (anonymous):

o whoops

OpenStudy (anonymous):

derivative of the denominator should be -1/x^2

OpenStudy (anonymous):

If you do that it keeps getting more complicated. I tried...

OpenStudy (anonymous):

the limit x->0- of e^1/x is 0

myininaya (myininaya):

that goes to infinity not zero kanade

OpenStudy (anonymous):

and theres a theorem that says that when lim x->a 1/0 = inf

OpenStudy (anonymous):

your coming in from the left not the right

OpenStudy (anonymous):

if you come in from the right it is infinity

OpenStudy (anonymous):

and so since u took l'hospitals once it is lim x->-0 -x^2/e^(1/x)

OpenStudy (anonymous):

but you need to remember that it is not exactly zero, it's approaching zero from the left

myininaya (myininaya):

ok you are right kadeo it is zero im sorry

OpenStudy (sriram):

i had earlier predicted that!!!

OpenStudy (anonymous):

im pretty sure u'll yeild different results if you approach the same equation from the right

OpenStudy (anonymous):

since the function isnt continuous

OpenStudy (anonymous):

Actually if you make it lim(x->0-) e^(-1/x) . ln(e^x) it gets a bit easier.

myininaya (myininaya):

why do you keep writing it as ln(e^x) instead of x? ln(e^x)=xlne=x(1)=x

myininaya (myininaya):

i just feel like im missing something or im misreading what you meant to say or said

OpenStudy (anonymous):

yea the ln(e^x) doesnt seem to hold much meaning...

OpenStudy (anonymous):

if it can just be written as x

OpenStudy (anonymous):

Yes, x e^(-1/x)

OpenStudy (anonymous):

OK it seems like nobody got it right, so here's the solution : lim (x->0-) lne^x/e^(1/x) = lim x . e(-1/x) = lim e(-1/x) / (1/x) = (De L'Hospital) = lim e(-1/x) . (1/x^2)/(-1/x^2) = - lim e (-1/x) = -inf